OpenStudy (goformit100):

A person who left home between 4 p.m. and 5 p.m. returned between 5 p.m. and 6 p.m. and found that the hands of his watch had exactly exchanged place, when did he go out ?

OpenStudy (goformit100):

@nathan917

OpenStudy (anonymous):

5:30 and 6:20?

OpenStudy (goformit100):

How ?

OpenStudy (anonymous):

do you have a grandfather clock at home?

OpenStudy (goformit100):

No My grand father is dead

OpenStudy (anonymous):

that is a grandfather clock

OpenStudy (goformit100):

ok

OpenStudy (agent0smith):

You should be able to find what you need here: https://sites.google.com/site/mymathclassroom/trigonometry/clock-angle-problems/clock-angle-problem-formula

OpenStudy (goformit100):

Ok Thank you Sir

OpenStudy (nathan917):

Yes. Thank you i was lost.

OpenStudy (anonymous):

left at 4:25 and came back at 5:20

OpenStudy (anonymous):

oops i was an hour off

OpenStudy (nathan917):

0.5˚ per minute and long hand moves 6˚ per minute let t₁ = time he left home in minutes. let t₂ = time he returned home in minutes. comparing the hour hand when he left and the minute hand when he returned 0.5*t₁ = 6*(t₂ - 300) t₁ = 12*(t₂ - 300) t₁ = 12t₂ - 3600 <---equation 1 comparing the hour hand when he returned and the minute hand when he left 0.5*t₂ = 6*(t₁ - 240) t₂ = 12*(t₁ - 240) t₂ = 12t₁ - 2880 <---equation 2 equating 1 and 2 will give us the ff. equation t₁ = 12(12t₁ - 2880) - 3600 t₁ = 144t₁ - 34560 - 3600 t₁ = 144t₁ - 38160 144t₁ - t₁ = 38160 143t₁ = 38160 t₁ = 266.8531469 or 4:26.85 pm t₂ = 12t₁ - 2880 t₂ = 12(266.8531469) - 2880 t₂ = 322.2377628 or 5:22.24 pm therefore he left the house at 4:26.85 pm and returned at 5:22.24 pm

OpenStudy (nathan917):

Got it.

OpenStudy (goformit100):

Thanks again